OpenStudy (anonymous):

Factor completely: 40xy^2 - 36x^2y. A. 4xy(9y - 10x) B. 4xy(10y - 9x) C. Prime D. 4xy(10x - 9y)

OpenStudy (campbell_st):

ok... whats th highest common factor of 40 and 36..? then whats the highest common factor of x and x^2 lastly what is the highest common factor of y^2 and y? these values will be outside the brackets

OpenStudy (anonymous):

4,X,Y

OpenStudy (campbell_st):

great so \[4 \times 10 \times x \times y \times y = 40xy\] and \[4 \times - 9 \times x \times x \times y = -36x^2y\] so if 4xy is outside the brackets what do you think is inside the brackets..? for each expanded expression just cross out 4, x and y

OpenStudy (anonymous):

10y-9x?

OpenStudy (campbell_st):

thats correct... so the answer is 4xy(10y - 9x) simple as that